# Self-Supervised Interpretable Sensorimotor Learning via Latent Functional Modularity

### v1

@@ Abstract
We introduce MoNet, a novel method that combines end-to-end learning with modular network architectures for self-supervised and interpretable sensorimotor learning. MoNet is composed of three functionally distinct neural modules: Perception, Planning, and Control. Leveraging its inherent modularity through a cognition-guided contrastive loss function, MoNet efficiently learns task-specific decision-making processes in latent space, without requiring task-level supervision. Moreover, our method incorporates an online post-hoc explainability approach, which enhances the interpretability of the end-to-end inferences without a trade-off in sensorimotor performance. In real-world indoor environments, MoNet demonstrates effective visual autonomous navigation, surpassing baseline models by 11% to 47% in task specificity analysis. We further delve into the interpretability of our network through the post-hoc analysis of perceptual saliency maps and latent decision vectors. This offers insights into the incorporation of explainable artificial intelligence within the realm of robotic learning, encompassing both perceptual and behavioral perspectives.
Introduction
One of the main objectives in end-to-end learning for autonomous navigation is to develop complex policies through human demonstrations. This is achieved by an end-to-end network that learns the hierarchical pipeline of perception, planning, and control in robotic systems via imitation learning (IL). Given that IL facilitates safe and efficient policy learning in an offline, supervised manner, end-to-end networks have been widely used in the design of learning-based applications (Tampuu et al. 2020).
However, although studies on IL have shown preliminary successes, designing an end-to-end sensorimotor network that can scale up to complex driving scenarios remains challenging. Traditional end-to-end networks often exhibit a less clear decision-making process, which complicates learning entangled tasks from demonstrations. To address this, recent conditional learning methods (Huang et al. 2020) employ multiple branching networks for each task, with outputs that switch based on task-level conditional inputs. However, this conditional input often corresponds to the outcome of an internal decision-making process in human demonstrations, which is typically implicit and difficult to identify. As a consequence, this approach necessitates extra task-level annotations (e.g., go-straight, turn-left), making it more demanding than simply collecting sensorimotor pairs.
Moreover, conventional networks, which directly compute control commands from sensory inputs, lack a transparent inference process. This obscurity makes it unclear what behavioral decision was intended for the resulting control output without direct execution. Several studies have sought to enhance interpretability by reconstructing various modalities (Chen, Li, and Tomizuka 2021; Zeng et al. 2019), or by visualizing attention maps (Kim and Canny 2017). However, they mainly focus on perceptual insights, still leaving the high-level decisions behind sensorimotor outputs largely obscure. This lack of clarity leads to insufficient task specificity and interpretability during the sensorimotor process, ultimately diminishing the reliability and trustworthiness of end-to-end networks in practical applications.
@@ Introduction
Figure 1: Our approach incorporates a functionally modular end-to-end network architecture, which includes a post-hoc process for an interpretable latent decision-making process.
To tackle those limitations, we adopt the insight of functional modularity in the neuroscience field for the design of an end-to-end architecture. Neuroscientists explain the organization of the human brain with two fundamental principles: functional segregation and integration (Meunier, Lambiotte, and Bullmore 2010). The brain has separate sub-modules that are structurally connected but functionally segregated according to their specific functions, such as sensory, cognitive, or motor processes, to perform complex behaviors. This suggests that employing a functionally modularized end-to-end architecture would be more suitable for approximating an autonomous driving stack comprising multiple sub-functions.
@@
In this paper, we present a modular end-to-end network, MoNet, inspired by the functional modularity of interconnected brain networks. MoNet is divided into perception, planning, and control modules, which are functionally separated but explicitly connected to form a single end-to-end structure (Fig. 1). Our network contains an internal latent decision process to facilitate task-oriented guidance for behaviorally relevant sensory-motor processes. Simultaneously, it employs a self-attention mechanism to extract salient spatial features from sensory input.
@@
Leveraging the inter-modular relationships in MoNet, we design a novel cognition-guided contrastive (CGC) loss function. Directed by the perception module with task-oriented contexts, this function encourages the planning module to make consistent decisions in similar driving contexts while differentiating responses in varied situations. Furthermore, we integrate a post-hoc multi-class classification method to decode task-relevant latent decisions into understandable representations. The internal hierarchy, combined with the contrastive scheme and augmented by the post-hoc explainability method, not only promotes functional specialization but also enables the emergence of a task-relevant and interpretable decision mechanism through self-supervision.
@@
In our evaluation, our method effectively demonstrates visual autonomous driving across multiple tasks, including corridor navigation, intersection navigation, and collision avoidance. We present empirical experiments conducted on a real-world robotic RC platform, showcasing the network’s capability to perform task-specific sensorimotor inference without requiring task-level labeling. We further explore spatial saliency maps and latent decisions during end-to-end inference. Specifically, by decoding latent decisions into explainable posterior probabilities, we gain the ability to visualize sequential high-level internal decisions alongside task uncertainty during continuous end-to-end sensorimotor control. These analyses highlight the significant interpretability and transparency of our end-to-end model, showcasing its effectiveness from both perceptual and behavioral perspectives in real-world continuous control applications.
@@
  • We propose MoNet, a modular end-to-end network that incorporates a post-hoc explainable method, enabling interpretable sensorimotor control.
  • We design a novel cognition-guided contrastive loss function to enhance the task-relevant and interpretable decision mechanism within the end-to-end network.
@@
In autonomous driving, end-to-end methods employ single neural networks to directly map sensory inputs to control outputs. ALVINN, the initial model for steering angle inference, utilized a multilayer perceptron (Pomerleau 1988). This approach has evolved to include convolutional neural networks (CNNs), mainly focused on lane-following tasks (Bojarski et al. 2016). Recent advancements have incorporated conditional imitation learning to cover a broader range of driving tasks (Gao et al. 2017; Codevilla et al. 2018; Huang et al. 2020; Zhang, Huang, and Ohn-Bar 2023). These methods use multiple branched layers switched by conditional inputs for navigating environments, such as ’go-straight’, ’turn-left’, or ’turn-right’. While such methods reduce task-level ambiguity, they necessitate additional human-engineered labeling for the navigational inputs and are constrained to predefined tasks. Moreover, interpreting the perceptual and behavioral processes within end-to-end networks remains a challenge, which affects confidence in the network’s reliability for real-world deployment.
@@
Recent studies have concentrated on designing interpretable end-to-end networks to address existing limitations. In this context, researchers using segmentation methods (Chen, Li, and Tomizuka 2021; Teng et al. 2022) have indirectly shown how a network can comprehend surrounding contexts by generating semantic masks from hidden features. Similarly, studies involving multi-head networks (Zeng et al. 2019) have evaluated the effectiveness of their planning methods by examining interpretable representations across various modalities, such as object detection or cost map generation. In contrast, attention mechanisms (Vaswani et al. 2017; Kim and Canny 2017) in recent studies have explicitly facilitated a deeper understanding of the areas within given feature elements where the network predominantly focuses during feedforward processing. Specifically, in the realm of autonomous driving, methods leveraging attention aim to accentuate critical aspects in driving scenarios, such as lane following (Shi et al. 2020), lane changing (Chen et al. 2019), or navigating intersections (Seong et al. 2021). However, the majority of research has primarily focused on the cognitive interpretations of how networks perceive contexts. Our work takes this a step further by investigating how to interpret the task-oriented intentions of the network in an explainable way. This approach enables both perceptual and behavioral interpretations online during end-to-end inference.
Figure 2: Overview of our method. While the entire end-to-end network is optimized by the supervised imitation loss $\mathcal{L}_{\pi}$, the planning module is updated by the cognition-guided contrastive loss $\mathcal{L}_{CGC}$, which is directed by the latent fusion vector $z^{p}$.
Modular End-to-End Network
Latent Functional Modularity
Our main idea is to embed functional modularity with internal hierarchy into an end-to-end network, allowing functionalities of the robotic sub-modules in latent space. As shown in Fig. 1, our modular end-to-end network, MoNet, has three distinct neural modules: Perception ($\mathcal{P}$), Planning ($\mathcal{Q}$), and Control $(\mathcal{R})$, which are the major components of the robotics system (Schwarting, Alonso-Mora, and Rus 2018). Each module 1) encodes raw observations $o$ into a fused perception feature vector $z^{p}$, 2) infers a latent decision $h^{d}$, and 3) computes a sensorimotor command ${a}^{c}$, respectively. The modules are functionally separated yet structurally connected in latent space, enabling them to constitute an end-to-end policy network $\pi$, parameterized by $\theta$:
@@
To encourage functional specialization of the modules in the network, we utilize two distinct mechanisms: bottom-up and top-down neural processes  (Baluch and Itti 2011; Anderson et al. 2018). Specifically, the bottom-up mechanism is a stimulus-driven, exogenous process, while the top-down mechanism is a behavior-relevant endogenous process (Katsuki and Constantinidis 2014). Considering their properties, the perception ($\mathcal{P}_{\theta}$) and planning ($\mathcal{Q}_{\theta}$) modules configure with self-attention mechanisms (Vaswani et al. 2017) to extract salient spatial features from sensory input $o$ and to obtain contextual importance from the features $z^{p}$, respectively. In contrast, the control ($\mathcal{R}_{\theta}$) module is designed with a top-down mechanism to internally modulate the sensory-motor signals based on the context-oriented behavioral decision $h^{d}$ from the planning module. This internal hierarchy enables the network to generate spatial attention maps and high-level latent decisions that are explicitly accessible during end-to-end inference. Employing a post-hoc approach allows these to be transformed into interpretable salient maps and behavioral decisions, respectively.
Network Details
@@
Our network receives a high-dimensional observation ${o}=[I,M]$ that includes a front camera image $I\in\mathbb{R}^{224\times 224\times 1}$ and a topology map $M\in\mathbb{R}^{64\times 64\times 3}$. This observation includes visual sensory data with navigational information, providing driving contexts in the ego-centric area for navigating complex environments, such as corridors with intersections. To effectively process high-dimensional camera images, we employ a hybrid architecture that combines the Vision Transformer encoder with CNN blocks (Dosovitskiy et al. 2020). In the perception module, the image $I$ and the topology map $M$ are first encoded into a hidden image feature $z_{I}\in\mathbb{R}^{6\times 6\times 64}$ and a hidden route feature $z_{M}\in\mathbb{R}^{1\times 1\times 64}$, respectively, using ResNet-inspired CNN blocks (He et al. 2016). The image feature is first reshaped into a flattened embedding $z^{\text{flat}}_{I}\in\mathbb{R}^{(6\times 6)\times 64}$, serving as a tokenized embedding for $N\!=6\!\times\!6$ image patches. This reshaped embedding is then concatenated with a positional embedding and fed into a Transformer encoder network (see the appendix for details). Subsequently, the Transformer model processes this input to produce an attention matrix $A(\cdot,\cdot)$, which integrates with the global context of the image feature $z_{I}$ via the self-attention mechanism. The attention matrix is computed as follows:
@@
where $Q,K,V\in\mathbb{R}^{N\times D_{k}}$ refer to queries, keys, and values consisting of $N$ data nodes with $D_{k}$ dimension size by following common terminology (Dosovitskiy et al. 2020). $W_{Q},W_{K},W_{V}\in\mathbb{R}^{D_{k}\times D}$ are weight matrices for an arbitrary input feature $Z\in\mathbb{R}^{N\times D}$ with $D$ dimension size to compute $Q,K,V$. The attention matrix $A\in\mathbb{R}^{N\times N}$ discerns the spatial significance of feature elements in the input image, offering valuable information to interpret the module’s bottom-up neural processing from a perceptual standpoint. Finally, by applying mean pooling, we derive the attention-integrated feature $z^{\text{att}}_{I}$ with reduced dimensionality. This feature vector is then concatenated with the flattened route feature $z^{\text{flat}}_{M}\in\mathbb{R}^{64}$, yielding a latent feature fusion vector $z^{p}=[z^{\text{att}}_{I},z^{\text{flat}}_{M}]$. The fused feature is subsequently fed into the planning and control modules without nonlinearity.
@@
This module extracts contextual features from the fused vector $z^{p}$ and produces a latent decision $h^{d}$ to modulate the neural signals of the control module in a top-down manner. We construct the planning module using another Transformer network, mirroring the encoder of the perception module. Here, the fusion vector is expanded and tokenized into an input embedding, corresponding to the embedding $z^{\text{flat}}_{I}$ in perception. This input embedding is then concatenated with a positional embedding, following a similar process to that in the perception module. To derive the latent decision in continuous space, we apply a linear layer to the output of the Transformer encoder without using a nonlinear activation function.
@@
Training Details
@@
where the loss term for throttle control $\tau^{c}$ is weighted by the parameter $\lambda_{\tau}\in[0,1]$. This weighting aims to emphasize supervision on steering control in visual autonomous navigation (Codevilla et al. 2018). Given that we collect noisy demonstrations from a real robot platform, we choose the L1 loss to reduce the penalty for large errors and be more robust to outliers compared to the L2 loss.
Furthermore, to enhance the distinctiveness of top-down latent decisions, we design a cognition-guided contrastive (CGC) loss function using a self-supervised approach, leveraging the modular characteristics of our end-to-end network. Generally, the output of planning is influenced by the context of the driving scene. This implies that similar situations lead to analogous decisions, while different scenarios result in distinct plans. Building on the observation that planning outputs are context-dependent, we define a cognition-guided contrastive loss, denoted as $\mathcal{L}_{\text{CGC}}({z}^{p}_{i},{z}^{p}_{j},{h}^{d}_{i},{h}^{d}_{j})$. Here, the latent decision $h^{d}$ is guided by the output feature fusion $z^{p}$ of the perception module as follows:
$\mathcal{L}_{\text{CGC}}=\begin{cases}\hfil 1-\text{cos}({h}^{d}_{i},{h}^{d}_{j})&\!\!\!\!\text{if}\ \text{cos}({z}^{p}_{i},{z}^{p}_{j})>=\kappa\\ \hfil\text{max}(0,\text{cos}({h}^{d}_{i},{h}^{d}_{j}))&\!\!\!\!\text{if}\ \text{cos}({z}^{p}_{i},{z}^{p}_{j})<\kappa\end{cases}$ (8)
where $\text{cos}(\alpha,\beta)=\frac{\alpha\cdot\beta}{|\alpha||\beta|}$ is cosine similarity that is widely used for similarity and clustering analysis in data science (Larose and Larose 2014). It calculates the distance between two vectors based on their relative orientations, rather than their absolute distance, within the bounded range [-1, +1]. The subscript $j$ represents the index of a sample within a mini-batch, different from the current sample index $i$. By minimizing Eq. 8, we aim to reduce the intra-cluster distance for latent decisions in demonstrations where perceptual feature similarity exceeds $\kappa$. Conversely, we strive to increase the inter-cluster distance when the similarity is below $\kappa$. This approach incentivizes the planning module to generate more consistent latent decisions for scenarios with comparable perceptual contexts, while ensuring diverse decisions for scenarios with differing contexts. Moreover, by leveraging perceptual features, our method eliminates the need to define positive or negative samples, thereby enabling contrastive learning through a self-supervised approach.
@@
$\mathcal{L}=\mathcal{L}_{\pi}({a^{c}_{i}},{a_{i}})+\lambda_{\text{CGC}}\mathcal{L}_{\text{CGC}}({z}^{p}_{i},{z}^{p}_{j},{h}^{d}_{i},{h}^{d}_{j})$ (9)
where $\lambda_{\text{CGC}}$ is a weight parameter. During the training phase, the supervised loss function $\mathcal{L}_{\pi}$ propagates the gradient flow across all modules ($\mathcal{P,Q,R}$), while the latent-guided loss function $\mathcal{L}_{\text{CGC}}$ targets only the planning and perception modules ($\mathcal{P,Q}$), promoting functional distinction between the planning and control modules ($\mathcal{Q,R}$).
Interpretation Details
Cognitive Interpretation:
We use the attention matrix of the perception module to create a saliency map $S$, which highlights the spatial regions in the current driving scene that the network focuses on from a cognitive perspective. The module generates the attention matrix $A\in\mathbb{R}^{N\times N}$ corresponding to the flattened vector of the encoded feature map $z_{I}\in\mathbb{R}^{h\times w\times c}$, where $N=h\times w$ is a resulting size of attention, $(h,w)$ is a reduced resolution of the image $I\in\mathbb{R}^{224\times 224\times 1}$, and $c$ is the feature dimension of $z_{I}$. Thus, we initially aggregate weights along the first dimension of $A$ to obtain the averaged attention weights $\bar{A}\in\mathbb{R}^{1\times N}$:
$\bar{A}_{j}=\frac{1}{N}\sum\nolimits_{i=1}^{N}A_{ij}\quad\textrm{for}\quad j=1,...N$ (10)

### v2

# Self-Supervised Interpretable End-to-End Learning via Latent Functional Modularity

@@ Abstract
We introduce MoNet, a novel functionally modular network for self-supervised and interpretable end-to-end learning. By leveraging its functional modularity with a latent-guided contrastive loss function, MoNet efficiently learns task-specific decision-making processes in latent space without requiring task-level supervision. Moreover, our method incorporates an online, post-hoc explainability approach that enhances the interpretability of end-to-end inferences without compromising sensorimotor control performance. In real-world indoor environments, MoNet demonstrates effective visual autonomous navigation, outperforming baseline models by 7% to 28% in task specificity analysis. We further explore the interpretability of our network through post-hoc analysis of perceptual saliency maps and latent decision vectors. This provides valuable insights into the incorporation of explainable artificial intelligence into robotic learning, encompassing both perceptual and behavioral perspectives. Supplementary materials are available at https://sites.google.com/view/monet-lgc.
End-to-End Learning, Self-Supervised Learning, Interpretable Learning, Modular Network, Imitation Learning
1 Introduction
@@
Figure 1: Our approach incorporates a functionally modular end-to-end network architecture, which includes a post-hoc method for an interpretable latent decision-making process.
One of the main objectives of end-to-end learning for autonomous navigation is to develop complex policies through human demonstrations. This is achieved by an end-to-end network that learns the hierarchical pipeline of perception, planning, and control in robotic systems via imitation learning (IL). Given that IL facilitates safe and efficient policy learning in an offline, supervised manner, end-to-end networks have been widely used in the design of learning-based applications (Tampuu et al., 2020).
However, although studies on IL have shown preliminary successes, designing an end-to-end sensorimotor network that can scale up to complex driving scenarios remains challenging. Traditional end-to-end networks often exhibit a less clear decision-making process, which complicates learning entangled tasks from demonstrations. To address this, recent conditional learning methods (Huang et al., 2020) employ multiple branching networks for each task, with outputs that switch based on task-level conditional inputs. However, this conditional input often corresponds to the outcome of an internal decision-making process in human demonstrations, which is typically implicit and difficult to identify. As a consequence, this approach necessitates extra task-level annotations (e.g., go-straight, turn-left), making it more demanding than simply collecting sensorimotor pairs.
Moreover, conventional networks, which directly compute control commands from sensory inputs, lack a transparent inference process. This obscurity makes it unclear what behavioral decision was intended for the resulting control output without direct execution. Several studies have sought to enhance interpretability by reconstructing various modalities (Chen et al., 2021; Zeng et al., 2019), or by visualizing attention maps (Kim & Canny, 2017). However, they mainly focus on perceptual insights, still leaving the high-level decisions behind sensorimotor outputs largely obscure. This lack of clarity leads to insufficient task specificity and interpretability during the sensorimotor process, ultimately diminishing the reliability and trustworthiness of end-to-end networks in practical applications.
@@
In this paper, we present a modular network, MoNet, a functionally modularized end-to-end architecture (Meunier et al., 2010) for autonomous navigation. MoNet is divided into perception, planning, and control modules, which are functionally separated but explicitly connected to form a single end-to-end structure (Fig. 1). Our network includes an internal latent decision process to facilitate task-oriented guidance for behaviorally relevant sensory-motor processes. Simultaneously, it employs a self-attention mechanism to extract salient spatial features from sensory input.
@@
Leveraging the modularity in MoNet, we design a novel self-supervised, latent-guided contrastive (LGC) loss function. Directed by latent features from the perception module with task-oriented contexts, this loss function encourages the planning module to make consistent decisions in similar driving contexts while differentiating responses in varied situations. The internal hierarchy, combined with our contrastive learning scheme, not only promotes functional specialization but also enables the emergence of a task-relevant decision-making mechanism through self-supervision.
Furthermore, we integrate a post-hoc technique from the field of explainable artificial intelligence (XAI) with our modular end-to-end network to transform task-relevant latent decisions into understandable representations. We implement a multi-class pattern classifier to predict the high-level task intent derived from these latent decisions. Subsequently, we calibrate the posterior probabilities of the classification results to achieve a more interpretable representation. These probabilities are then converted into an entropy value, which quantifies the uncertainty of the end-to-end model’s inference from a task-level perspective.
@@
In our evaluation, our method effectively demonstrates visual autonomous driving across multiple tasks, including corridor navigation, intersection navigation, and collision avoidance. We present empirical experiments conducted on a real-world robotic RC platform, showcasing the network’s capability to perform task-specific sensorimotor inference without requiring task-level labeling. We further explore spatial saliency maps and latent decisions during end-to-end navigation in the real world. Specifically, by decoding latent decisions into explainable posterior probabilities, we gain the ability to visualize sequential high-level internal decisions alongside task uncertainty during continuous end-to-end sensorimotor control. These analyses highlight the significant interpretability and transparency of our end-to-end model, showcasing its effectiveness from both perceptual and behavioral perspectives in real-world continuous control applications.
@@
  • We propose MoNet, a modular end-to-end network that incorporates a post-hoc explainability method, enabling interpretable sensorimotor control.
  • We design a self-supervised, latent-guided contrastive learning scheme to enhance the task-relevant decision-making mechanism within the end-to-end architecture.
@@ 2 Related Works
In autonomous driving, end-to-end methods employ single neural networks to directly map sensory inputs to control outputs. ALVINN, the initial model for steering angle inference, utilized a multilayer perceptron (Pomerleau, 1988). This approach has evolved to include convolutional neural networks (CNNs), mainly focused on lane-following tasks (Bojarski et al., 2016). Recent advancements have incorporated conditional imitation learning to cover a broader range of driving tasks (Gao et al., 2017; Codevilla et al., 2018; Huang et al., 2020; Zhang et al., 2023). These methods use multiple branched layers switched by conditional inputs for navigating environments, such as ’go-straight’, ’turn-left’, or ’turn-right’. While such methods reduce task-level ambiguity, they necessitate additional human-engineered labeling for the navigational inputs and are constrained to predefined tasks. Moreover, interpreting the perceptual and behavioral processes within end-to-end networks remains a challenge, which affects confidence in the network’s reliability for real-world deployment.
Figure 2: Overview of our method. While the entire end-to-end network is optimized by the supervised imitation loss $\mathcal{L}_{\pi}$, the planning module is updated by the latent-guided contrastive loss $\mathcal{L}_{LGC}$, which is directed by the latent vector $z^{p}$.
@@
Recent studies have concentrated on designing interpretable end-to-end networks to address existing limitations. In this context, researchers using segmentation methods (Chen et al., 2021; Teng et al., 2022) have indirectly shown how a network can comprehend surrounding contexts by generating semantic masks from hidden features. Similarly, studies involving multi-head networks (Zeng et al., 2019) have evaluated the effectiveness of their planning methods by examining interpretable representations across various modalities, such as object detection or cost map generation. In contrast, attention mechanisms (Vaswani et al., 2017; Kim & Canny, 2017) in recent studies have explicitly facilitated a deeper understanding of the areas within given feature elements where the network predominantly focuses during feedforward processing. Specifically, in the realm of autonomous driving, methods leveraging attention aim to accentuate critical aspects in driving scenarios, such as lane following (Shi et al., 2020), lane changing (Chen et al., 2019), or navigating intersections (Seong et al., 2021). However, the majority of research has primarily focused on the cognitive interpretations of how networks perceive contexts. Our work takes this a step further by investigating how to interpret the task-oriented intentions of the network in an explainable way. This approach enables both perceptual and behavioral interpretations online during end-to-end inference.
3 Modular End-to-End Network
3.1 Latent Functional Modularity
Our main idea is to embed functional modularity with internal hierarchy into an end-to-end network, allowing functionalities of the robotic sub-modules in latent space. As shown in Fig. 1, our modular end-to-end network, MoNet, has three distinct neural modules: Perception ($\mathcal{P}$), Planning ($\mathcal{Q}$), and Control $(\mathcal{R})$, which are the major components of the robotics system (Schwarting et al., 2018). Each module 1) encodes raw observations $o$ into a fused perception feature vector $z^{p}$, 2) infers a latent decision $h^{d}$, and 3) computes a sensorimotor command ${a}^{c}$, respectively. The modules are functionally separated yet structurally connected in latent space, enabling them to constitute an end-to-end policy network $\pi$, parameterized by $\theta$:
@@ 3.1 Latent Functional Modularity
To encourage functional specialization of the modules in the network, we utilize two distinct mechanisms: bottom-up and top-down neural processes  (Baluch & Itti, 2011; Anderson et al., 2018). Specifically, the bottom-up mechanism is a stimulus-driven, exogenous process, while the top-down mechanism is a behavior-relevant endogenous process (Katsuki & Constantinidis, 2014). Considering their properties, the perception ($\mathcal{P}_{\theta}$) and planning ($\mathcal{Q}_{\theta}$) modules configure with self-attention mechanisms (Vaswani et al., 2017) to extract salient spatial features from sensory input $o$ and to obtain contextual importance from the features $z^{p}$, respectively. In contrast, the control ($\mathcal{R}_{\theta}$) module is designed with a top-down mechanism to internally modulate the sensory-motor signals based on the context-oriented behavioral decision $h^{d}$ from the planning module. This internal hierarchy enables the network to generate spatial attention maps and high-level latent decisions that are explicitly accessible during end-to-end inference. Employing a post-hoc approach allows these to be transformed into interpretable salient maps and behavioral decisions, respectively.
3.2 Network Details
@@
Our network receives a high-dimensional observation ${o}=[I,M]$ that includes a front camera image $I\in\mathbb{R}^{224\times 224\times 1}$ and a topology map $M\in\mathbb{R}^{64\times 64\times 3}$. This observation includes visual sensory data with navigational information, providing driving contexts in the ego-centric area for navigating complex environments, such as corridors with intersections. To effectively process high-dimensional camera images, we employ a hybrid architecture that combines the Vision Transformer encoder with CNN blocks (Dosovitskiy et al., 2020). In the perception module, the image $I$ and the topology map $M$ are first encoded into a hidden image feature $z_{I}\in\mathbb{R}^{6\times 6\times 64}$ and a hidden route feature $z_{M}\in\mathbb{R}^{1\times 1\times 64}$, respectively, using ResNet-inspired CNN blocks (He et al., 2016). The image feature is first reshaped into a flattened embedding $z^{\text{flat}}_{I}\in\mathbb{R}^{(6\times 6)\times 64}$, serving as a tokenized embedding for $N\!=6\!\times\!6$ image patches. This reshaped embedding is then concatenated with a positional embedding and fed into a Transformer encoder network (see Appendix B.1). Subsequently, the Transformer model processes this input to produce an attention matrix $A(\cdot,\cdot)$, which integrates with the global context of the image feature $z_{I}$ via the self-attention mechanism. The attention matrix is computed as follows:
@@
where $Q,K,V\in\mathbb{R}^{N\times D_{k}}$ refer to queries, keys, and values consisting of $N$ data nodes with $D_{k}$ dimension size by following common terminology (Dosovitskiy et al., 2020). $W_{Q},W_{K},W_{V}\in\mathbb{R}^{D_{k}\times D}$ are weight matrices for an arbitrary input feature $Z\in\mathbb{R}^{N\times D}$ with $D$ dimension size to compute $Q,K,V$. The attention matrix $A\in\mathbb{R}^{N\times N}$ discerns the spatial significance of feature elements in the input image, offering valuable information to interpret the module’s bottom-up neural processing from a perceptual standpoint. Finally, by applying mean pooling, we derive the attention-integrated feature $z^{\text{att}}_{I}\in\mathbb{R}^{65}$ with reduced dimensionality. This feature is then concatenated with the flattened route feature $z^{\text{flat}}_{M}\in\mathbb{R}^{64}$, yielding a latent feature fusion vector $z^{p}=[z^{\text{att}}_{I},z^{\text{flat}}_{M}]$. The fused feature is then fed into the planning and control modules without nonlinearity.
@@
This module extracts contextual features from the fused vector $z^{p}$ and produces a latent decision $h^{d}$ to modulate the neural signals of the control module in a top-down manner. We construct the planning module using another Transformer network, mirroring the encoder of the perception module. Here, the fusion vector is expanded and tokenized into an input embedding $z^{emb}\in\mathbb{R}^{(65+64)\!\times\!64}$, corresponding to the embedding $z^{\text{flat}}_{I}$ in perception. This input embedding is then concatenated with a positional embedding, following the same process as that in the perception module. To derive the latent decision in continuous space, we apply a linear layer to the output of the Transformer encoder without using a nonlinear activation function.
@@
Figure 3: Our self-supervised contrastive learning scheme assesses the similarity of the perceptual features to decide on positive and negative latent decision samples.
3.3 Training Details
@@
where the loss term for throttle control $\tau^{c}$ is weighted by the parameter $\lambda_{\tau}\in[0,1]$. This weighting aims to emphasize supervision on steering control in visual autonomous navigation (Codevilla et al., 2018). Given that we collect noisy demonstrations from a real robot platform, we choose the L1 loss to reduce the penalty for large errors and be more robust to outliers compared to the L2 loss.
Figure 4: Overview of our post-hoc behavior interpretation process.
Furthermore, to enhance the distinctiveness of top-down latent decisions, we design a latent-guided contrastive (LGC) loss function using a self-supervised approach, leveraging the modular characteristics of our end-to-end network (Fig. 3). Generally, the output of planning is influenced by the context of the driving scene. This implies that similar situations lead to analogous decisions, while different scenarios result in distinct plans. Building on the observation that planning outputs are context-dependent, we define the latent-guided contrastive loss, denoted as $\mathcal{L}_{\text{LGC}}({z}^{p}_{i},{z}^{p}_{j},{h}^{d}_{i},{h}^{d}_{j})$. Here, the latent decision $h^{d}$ is guided by the feature fusion $z^{p}$ of the perception module as follows:
$\mathcal{L}_{\text{LGC}}=\begin{cases}\hfil 1-\text{cos}({h}^{d}_{i},{h}^{d}_{j})&\!\!\!\!\text{if}\ \text{cos}({z}^{p}_{i},{z}^{p}_{j})>=\kappa\\ \hfil\text{max}(0,\text{cos}({h}^{d}_{i},{h}^{d}_{j}))&\!\!\!\!\text{if}\ \text{cos}({z}^{p}_{i},{z}^{p}_{j})<\kappa\end{cases}$ (8)
where $\text{cos}(\alpha,\beta)=\frac{\alpha\cdot\beta}{|\alpha||\beta|}$ is cosine similarity that is widely used for similarity and clustering analysis in data science (Larose & Larose, 2014). It calculates the distance between two vectors based on their relative orientations, rather than their absolute distance, within the bounded range [-1, +1]. The subscript $j$ represents the index of a sample within a mini-batch, different from the current sample index $i$. By minimizing Eq. 8, we aim to reduce the intra-cluster distance for latent decisions in demonstrations where perceptual feature similarity exceeds $\kappa$. Conversely, we strive to increase the inter-cluster distance when the similarity is below $\kappa$. This approach incentivizes the planning module to generate more consistent latent decisions for scenarios with comparable perceptual contexts, while ensuring diverse decisions for scenarios with differing contexts. Moreover, by leveraging perceptual features, our method eliminates the need to define positive or negative samples, thereby enabling contrastive learning through a self-supervised approach.
@@
$\mathcal{L}=\mathcal{L}_{\pi}({a^{c}_{i}},{a_{i}})+\lambda_{\text{LGC}}\mathcal{L}_{\text{LGC}}({z}^{p}_{i},{z}^{p}_{j},{h}^{d}_{i},{h}^{d}_{j})$ (9)
where $\lambda_{\text{LGC}}$ is a weight parameter. During the training phase, the supervised loss function $\mathcal{L}_{\pi}$ propagates the gradient flow across all modules ($\mathcal{P,Q,R}$), while the latent-guided loss function $\mathcal{L}_{\text{LGC}}$ targets only the planning and perception modules ($\mathcal{P,Q}$), promoting functional distinction between the planning and control modules ($\mathcal{Q,R}$).
3.4 Interpretation Details
Perceptual Interpretation:
We use the attention matrix of the perception module to create a saliency map $S$, which highlights the spatial regions in the current driving scene that the network focuses on from a perceptual perspective. The module generates the attention matrix $A\in\mathbb{R}^{N\times N}$ corresponding to the flattened vector of the encoded feature map $z_{I}\in\mathbb{R}^{h\times w\times c}$, where $N=h\times w$ is a resulting size of attention, $(h,w)$ is a reduced resolution of the image $I\in\mathbb{R}^{224\times 224\times 1}$, and $c$ is the feature dimension of $z_{I}$. Thus, we initially aggregate weights along the first dimension of $A$ to obtain the averaged attention weights $\bar{A}\in\mathbb{R}^{1\times N}$:
$\bar{A}_{j}=\frac{1}{N}\sum\nolimits_{i=1}^{N}A_{ij}\quad\textrm{for}\quad j=1,...,N$ (10)